OpenStudy (anonymous):

How many grams of Fe can be produced when 5.50g of Fe2O3 reacts?

OpenStudy (somy):

decomposition reaction? or Fe2O3 reacts with something?

OpenStudy (somy):

i think you can do it pretty mathematically

OpenStudy (somy):

5.50g - Molar mass of Fe2O3 X - Molar mass of 2 Fe atoms

OpenStudy (somy):

@thomaster do u think its right?

OpenStudy (anonymous):

Is there more to the question?

OpenStudy (somy):

i think its right im pretty sure

OpenStudy (somy):

@Kainui what do u think?

OpenStudy (anonymous):

\[5.50g \times \frac{ 1 mol }{ 159.7 g} \times \frac{ 2 mol }{ 1mol} \times \frac{ 55.85 g }{ 1 mol} \]

OpenStudy (anonymous):

Like that @sonny

OpenStudy (anonymous):

I meant that as a question

OpenStudy (somy):

yup the answer is same its just i used mathematical way and you used formula

OpenStudy (anonymous):

I'm not sure if I understand what the question is asking

OpenStudy (anonymous):

I see

OpenStudy (somy):

its simply asking for mass of Fe in the compound

OpenStudy (anonymous):

Ok

OpenStudy (kainui):

Looks like you know what you're talking about and how to explain it well @Somy =D

OpenStudy (somy):

thanx :* @Kainui ❤_❤